OpenStudy (anonymous):

Which set of points lies on the graph of x = -2? (0, -2), (2, -2), (-2, -2) (-2, 0), (-2, 2), (-2, -2) (-2, 0), (2, -4), (-4, 2)

OpenStudy (anonymous):

(-2, 0), (-2, 2), (-2, -2)

OpenStudy (anonymous):

second set

OpenStudy (anonymous):

If y = 2x - 3, then which of the following ordered pairs lies on the graph? (1, -1) (-3, 0) (5, 4)

OpenStudy (anonymous):

first

OpenStudy (anonymous):

(1,-1)

OpenStudy (anonymous):

want me to work it out for you

OpenStudy (anonymous):

Yes, please! & if you can help me with others please message me!

OpenStudy (anonymous):

if y=2x-3 then you plug the number you have into the problem so for the first one y=2(1)-3 2x1=2 then 2-3=-1 so y=-1

OpenStudy (anonymous):

Which of the following points does not lie on the graph of y = 5x - 5? (0, -5) (1, 0) (-5,0)

OpenStudy (anonymous):

(1, 0)

OpenStudy (anonymous):

If the ordered pair (-5/2, y) lies on the graph of y = 7 - 2x, find the value of y. 2 12 19/2

OpenStudy (anonymous):

12

OpenStudy (anonymous):